OpenStudy (kuoministers):

Another factorising question (x^4 - 2)(x^4 - 1) Help pls :(

OpenStudy (zehanz):

You can use the form\[a^2-b^2=(a+b)(a-b)\]\[x^4=(x^2)^2\]so\[(x^4-2)(x^4-1)=(x^2+\sqrt{2})(x^2-\sqrt{2})(x^2+1)(x^2-1)\]You can do this again, because the second and fourth factor are again the difference of two squares.\[(x^2+\sqrt{2})(x+\sqrt[4]{2})(x-\sqrt[4]{2})(x^2+1)(x+1)(x-1)\] Now if you know what complex numbers are, you could even factorise the \[a^2+b^2\]factors, but for now this should do ;) ZeHanz

OpenStudy (kuoministers):

Wow thanks

OpenStudy (zehanz):

You're welcome!